# Cosmological Zoom-In Simulation of Odd Radio Circles as Merger-Driven Shocks in Galaxy Groups

### v1

@@ 3.6 Shock indicators
The built-in shock finder in the code OpenGadget3 is based on a geometrical on-the-fly scheme, that utilizes pressure gradients to construct the shock normal, along which then the Rankine-Hugoniot jump conditions are evaluated. After masking cold flows and shear flows based on threshold pressure and velocity jumps, the algorithm is finally inferring shock quantities such as the compression ratio and Mach number from approximated hydrodynamic quantities up- and downstream of the shock (Beck et al. 2016a). As mentioned in Sect. 3.5, the accuracy of the shock finder is strongly dependent on the resolution of the simulation. Specifically, the size of the local smoothing kernel of the respective SPH particle is defining the scale on which shocks are resolved. In our case, the mean smoothing scale of the simulated ORC is on the order of at , i.e. when the size and morphology matches best to observed counterparts. For reference, this is twice as big as the smoothing kernel imitating the resolution limit of ASKAP at , indicated by the small circle in the lower left corner of the polarization map in Fig. 5. This is the reason for the appearance of the spot-like features in the inferred mock radio images – in particular in the polarization features – which are encompassing the extent of the SPH particles in these regions.
@@
In order to illustrate the actual extent and width of the shock region, it is instructive to compare the results with an alternative way to determine the shock position. In particular, gaussian filtering of pressure maps has proven to be quite efficient in revealing pressure jumps that indicate a shock in the medium (see also Dolag et al. 2005b; Ragagnin et al. 2017; Sommer et al. 2024). Naturally, this approach is only feasible during post processing of the simulation, since it requires a 2D projection in the correct plane and visual inspection. Nevertheless it is a convenient tool to double check the conclusion of the shock finder. The process is the following: starting from a pressure map , we smooth the image by applying a 2D gaussian kernel function on each pixel of the original image and obtain a blurred variant . Sharp edges in the original map will still be present, but are naturally attenuated in the new image. If we now subtract the two images from each other, features spanning over a larger scale than the smoothing kernel will vanish and only sharp features remain. Lastly, the contribution of each pixel is normalized by the local value of the blurred image, simply to see each edge feature in the resulting map equally strong, no matter the pressure contrast. The final shock map will be therefore given by
@@
In Fig. 6 we present such a shock map of the system which was generated by the method explained above. The time and scale of the simulated ORC is again the same as in the polarization map in Fig. 5. The upper panel is showing the integrated surface density of the gas, while the bottom panel is displaying a shock map in the same scale. The pressure at each pixel was determined via the Compton- parameter, which can be used as a proxy for the integrated electron pressure along the line of sight. Finally, we used a gaussian smoothing kernel length of pixels, which is corresponding to approximately in this case. The shock map in Fig. 6 is revealing the clear edges in pressure, displaying its strikingly circular geometry which was not as clear in the output of the shock finder.
@@ 4 Discussion
We have shown in Sect. 3, that merger-accelerated shocks can produce giant circular radio features around massive galaxies and galaxy groups. Their morphology closely resembles those of a recently discovered class of radio objects, so called Odd Radio Circles, also dubbed ORCs (e.g. Norris et al. 2021, 2022; Koribalski et al. 2021, 2024b). It is yet under debate as to what these peculiar structures are, since they have a number of unusual traits. Particularly, they display i) a remarkably circular, edge-brightened morphology, ii) have large sizes comparable to the virial radius of a central elliptical galaxy group and iii) exhibit exceedingly high radio powers, judging by self-similar scaling of their halo mass (c.f. Figure 6 by Koribalski et al. 2024b).
@@
The effectiveness of galaxy mergers in producing circular radio features matching ORCs was first shown in a pilot study by Dolag et al. (2023), where the authors have conducted cosmological zoom-in simulations of a Milky-Way-like halo with a mass of approximately . Since the simulations were performed in a pure hydrodynamics fashion without magnetic fields, they needed to estimate the radio brightness during postprocessing. That was realized by adding CRs to shock regions, where their amount was chosen such that the integrated CR energy was a fixed fraction of the dissipated shock energy, based on the shock finder output at a single timestep. Finally by assuming a constant and homogeneous magnetic field, they could infer the spectral radio flux based on synchrotron emission. Although the morphology and size of ORCs could be reproduced, these simulations predicted radio powers that were far too low compared to what is measured in ORCs. However it was unclear which assumptions specifically have led to this deficiency. In this work, we are closing two main shortcomings of the study, namely by simulating i) a massive enough galaxy group (), comparable to what is found in observed counterparts and ii) using a magneto-hydrodynamic scheme with self-consistent, on-the-fly treatment of spectral CRs.
@@
As discussed in Sect. 3.3, it is necessary to proceed with caution when evaluating the absolute value of the magnetic field in a cosmological simulation. It tends to be underestimated, because technically one needs extremely high resolution in order to resolve the turbulent dynamo process in the intergalactic medium, which is the main driver in amplifying the magnetic field (e.g. Vazza et al. 2014, 2018; Böss et al. 2024; Steinwandel et al. 2024). Since the radio properties of the system are based on synchrotron emission, this can significantly alter the final result. To obtain more realistic predictions, it is therefore appropriate to compare the simulated magnetic field strength with analytic predictions for the system. In Fig. 7, we show the time-dependent radio luminosity due to synchrotron emission at for the simulated ORC. The three vertical lines represent the three times from the visualization in Fig. 1, which are each about apart, starting from shortly before the merger. The colored regions represent three different models for the magnetic field strength, where the lower end is the prediction by the respective model and the upper end is representing a boost by a factor of 10. The orange area is the luminosity profile determined by the simulated magnetic field strength , while the other two were generated by assuming a balance between i) magnetic field pressure and thermal pressure (, blue) and ii) magnetic field pressure and turbulent pressure (, red). Hence the local magnetic field strength is given by
@@
in the two cases respectively, where is the thermal pressure, the mass density and the turbulent velocity. It is defined as , where is the the local velocity, while the local bulk velocity is computed as the kernel weighted mean velocity of the neighboring particles
@@
In order to keep the models overall realistic, the value of the plasma- and the balance factor are scaled for the whole simulation, such that the field strength in the central regions of the merging halos is in the order of several , which is in accordance to what is measured at galaxies in that mass range (e.g. Chyży & Beck 2004; Opher et al. 2009; Nikiel-Wroczyński et al. 2020). As expected, the two models generally produce larger values for the total magnetic field strength than the simulated one, reflected by the systematically larger inferred radio power. The turbulent model is overall the most powerful one, which is particularly visible at later times.
@@
The results from the pilot study by Dolag et al. (2023) introduced in the beginning of this Section are visualized as the horizontal grey band in Fig. 7, where its width is representing the overall extent of the total radio power inferred in postprocessing of their simulation. Compared to their estimate, our simulation is exceeding it by at least three orders of magnitude at moment of peak brightness – already when using the radio power inferred from the simulated magnetic field strength alone. But even taking the boost of stronger field models into account, we are still not reaching the luminosity of observed counterparts indicated by the various dashed-dotted lines in Fig. 7. Consequently, also the surface brightness is underpredicted by the simulation. The ORC Physalis, for example, has a flux density in the order of in the outer radio rings (c.f. bottom panel of Fig. 7), while the simulated counterpart is about three orders of magnitude lower at (cf. Fig. 3). This means that, although we are now resolving a halo in the right mass regime and reproduce the correct X-ray properties (c.f. Sect. 3.2), there is clearly still an aspect in the picture missing, in order to entirely explain the physical nature of these systems.
@@ 4.2 Potential role of AGN and stellar feedback
Assuming that there is no other essential process producing significant radio emission other than synchrotron emission, this lack of total radio luminosity would then mean that either there are not enough relativistic CR electrons present, or/and the local magnetic field strength is too low. Since the maximum Mach number of the shock, when the simulated ORC is also the brightest, is already relatively large (, c.f. Sect. 3.4), a stronger shock would not remedy the CR deficiency, because their acceleration efficiency via DSA would not increase substantially. A stronger magnetic field would in principle also give rise to stronger synchrotron emission. The problem is, that several processes are currently underrepresented by our magnetic field models, which are presumed to contribute to the state of the magnetic field in the medium, e.g. Bell amplification in shocks or magnetization of the halo by highly magnetized or laminar outflows. To test the effect of stronger magnetic fields in our simulation, we would need to increase the field strength across the whole system in order to stay self consistent. But this would push the field strength in the central regions of the galaxies beyond actually observed values, and is therefore not an option. Additionally, we need to consider the relative importance of fossil CR populations in the halo, which would be present in case of prior injection events of CRs into the CGM of the merging galaxies. Typically, enough time would have passed such that these old populations would have cooled down to mildly relativistic energies – meaning that they would not be radio-bright anymore – but still not enough time for those populations to join the thermal pool of electrons. While a strong shock is necessary to accelerate charged particles from the Maxwellian distribution, a relatively weak shock would already suffice to push such fossil CRs back to relativistic energies again, effectively boosting the total radio brightness of the whole system (e.g. Caprioli et al. 2018; Kang 2018).
@@
Apart from merger or accretion shocks, stellar outflows and AGN feedback are the other main sources for these fossil populations and it can be assumed that they should play a role in the two galaxies in our simulation. But since it was performed in a non-radiative fashion, meaning that it did not include star formation, it is not possible to make definite statements about the actual contribution of stars and black holes at this point. A reasonable assumption would be that they would raise the amount of relativistic electrons by a factor of 10, which is reflected by the upper end of the inferred luminosity profiles of the simulation in Fig. 7. But evidently, that would not suffice to reach the observed values. Nevertheless, the fractional polarization is actually also pointing towards the importance of fossil CRs in this system. The simulation produces values mainly between (c.f. Sect. 3.5), while measurements of ORC 1 report lower values mostly between in the interior and reaching 30% at most in the ring (Norris et al. 2022). If there are significant amounts of old CR populations present, one would certainly increase the number of emission centers along the line of sight compared to the case where electrons were accelerated from the thermal pool only. But since also the direction of the magnetic field vector is varying along the line of sight – and the polarization angle is depending on that direction – this would naturally diversify the distribution of polarization angles present in the total emission. Therefore, the total polarization fraction along a given line of sight would decrease in the case of additional fossil CRs populations, which is necessary to get into agreement with observations. Thus, in order to give a concluding answer to the question whether fossil CR populations are actually the missing key to explain all radio features of ORCs, one would need to perform the simulation again with star formation and proper feedback implementation including self-consistent CR treatment, which is a challenge to current state-of-the-art cosmological codes and we will refer to in future work.
@@ 5 Summary and conclusion
In this article, we have presented a first of its kind zoom-in simulation of a galaxy group performed in a magnetohydronamic setup with spectral cosmic ray treatment on-the-fly. The scientific aim was to test the limits of a strong contender for a formation channel for Odd Radio Circles (ORCs) – a new class of radio objects whose origin is still a puzzle. The idea is that during a major merger event between two galaxies, the immense energy dissipation can release several strong bow shocks (i.e. merger shocks). If the projection is close to the perpendicular direction to the merger axis, these shocks can form a giant ring, which closely resembles the edge-brightened radio features that are now identified as ORCs in observations.
This simulation incorporates two major improvements compared to the pilot study by Dolag et al. (2023), who introduced this formation path: it is i) resolving a 10 times more massive halo with a final virial mass of about , which is now in agreement with the group mass measured in the center of ORCs and ii) determining CR population and local magnetic field direction self-consistently, rather than adding them in postprocessing. We showed that the simulation reproduces the X-ray surface brightness of a galaxy group in that mass range and demonstrated that the peculiar offset between the X-ray and radio-bright centers observed in ORC Physalis (Koribalski et al. 2024b) can be explained by the difference in gas density at the centers of the merging galaxies (Sect. 3.2). By providing mock observations of the radio surface brightness and spectral index based on working frequencies of LOFAR, GMRT and ASKAP, we demonstrated that the morphology and size agrees well with observations (Sect. 3.3). The emission spectrum of the radio shells is indicating a sonic Mach number of , which is comparable to strong shock events in massive galaxy clusters (Sect. 3.4). The simulation produces a fractional polarization degree of across the radio ring (Sect. 3.5). The pressure excess map provided in Sect. 3.6 reveals the sharp shock boundaries responsible for the radio features.
Compared to observations, the simulation is underpredicting the total radio luminosity and overestimating its polarization degree (Sect. 4). Both issues would be alleviated by factoring in additional sources for relativistic cosmic rays apart from DSA, which is the only acceleration mechanism in the present simulation. Such mechanisms would be the injection of cosmic rays via stellar or AGN feedback, which could significantly contribute to the total energy budget and boost the radio brightness when reaccelerated by the merger shock. We will implement a description of these fossil cosmic rays in our future work and investigate whether they could be the missing ingredient for explaining the observed radio power of ORCs.

### v2

@@ 3.6 Shock indicators
The built-in shock finder in the code OpenGadget3 is based on a geometrical on-the-fly scheme that utilises pressure gradients to construct the shock normal along which the Rankine-Hugoniot jump conditions are evaluated. After masking cold flows and shear flows based on threshold pressure and velocity jumps, the algorithm finally inferes shock quantities such as the compression ratio and Mach number from approximated hydrodynamic quantities up- and downstream of the shock (Beck et al. 2016a). As mentioned in Sect. 3.5, the accuracy of the shock finder strongly depends on the resolution of the simulation. Specifically, the size of the local smoothing kernel of the respective SPH particle defines the scale on which the shocks are resolved. In our case, the mean smoothing scale of the simulated ORC is of the order of at , i.e. when the size and morphology matches best to observed counterparts. For reference, this is twice as big as the smoothing kernel imitating the resolution limit of ASKAP at , indicated by the small circle in the lower left corner of the polarisation map in Fig. 5. This explains the appearance of the spot-like features in the inferred mock radio images – particularly in the polarisation features – which encompass the extent of the SPH particles in these regions.
@@
In order to illustrate the actual extent and width of the shock region, it is instructive to compare the results with an alternative way to determine the shock position. In particular, Gaussian filtering of pressure maps has proven to be quite efficient in revealing pressure jumps that indicate a shock in the medium (see also Dolag et al. 2005b; Ragagnin et al. 2017; Sommer et al. 2024). Naturally, this approach is only feasible during post-processing of the simulation, since it requires a 2D projection in the correct plane and visual inspection. Nevertheless, it is a convenient tool for double-checking the conclusion of the shock finder. The process is as follows: starting from a pressure map , we smoothed the image by applying a 2D Gaussian kernel function to each pixel of the original image, yielding a blurred variant . Sharp edges in the original map remain present, but are naturally attenuated in the new image. If we now subtract the two images from each other, features spanning over a larger scale than the smoothing kernel vanish and only sharp features remain. Lastly, the contribution of each pixel was normalised by the local value of the blurred image, to ensure all edge features appear equally strong in the resulting map, no matter the pressure contrast. The final shock map is therefore given by
@@
In Fig. 6 we present such a shock map of the system generated by the method explained above. The time and scale of the simulated ORC is again the same as in the polarisation map in Fig. 5. The upper panel shows the integrated surface density of the gas, while the bottom panel displays a shock map in the same scale. The pressure at each pixel was determined via the Compton- parameter, which can be used as a proxy for the integrated electron pressure along the line of sight. Finally, we used a Gaussian smoothing kernel length of pixels, which corresponds to approximately in this case. The shock map in Fig. 6 reveals clear pressure edges, with strikingly circular geometry, which was not evident in the shock finder output.
@@ 4 Discussion
We showed in Sect. 3, that merger-accelerated shocks can produce giant circular radio features around massive galaxies and galaxy groups. Their morphology closely resembles those of a recently discovered class of radio objects, so-called ORCs (e.g. Norris et al. 2021b, 2022; Koribalski et al. 2021, 2024b). It is yet under debate as to what these peculiar structures are, since they have a number of unusual traits. Particularly, they display i) a remarkably circular, edge-brightened morphology, ii) have large sizes comparable to the virial radius of a central elliptical galaxy group and iii) exhibit exceedingly high radio powers, judging by self-similar scaling of their halo mass (cf. Figure 6 by Koribalski et al. 2024b).
@@
The effectiveness of galaxy mergers in producing circular radio features matching ORCs was first shown in a pilot study by Dolag et al. (2023), where the authors conducted cosmological zoom-in simulations of a Milky-Way-like halo with a mass of approximately . Since the simulations were performed in a pure hydrodynamics fashion without magnetic fields, they needed to estimate the radio brightness during post-processing. That was realised by adding CRs to shock regions, where their amount was chosen such that the integrated CR energy was a fixed fraction of the dissipated shock energy, based on the shock finder output at a single time step. Finally, by assuming a constant and homogeneous magnetic field, they could infer the spectral radio flux based on synchrotron emission. Although the morphology and size of ORCs could be reproduced, these simulations predicted radio powers that were far too low compared to what is measured in ORCs. However, it was unclear which assumptions specifically led to this deficiency. In this work, we address two main shortcomings of the study by simulating i) a massive galaxy group (), comparable to observed counterparts and ii) using a magnetohydrodynamic scheme with a self-consistent, on-the-fly treatment of spectral CRs.
@@
As discussed in Sect. 3.3, it is necessary to proceed with caution when evaluating the absolute value of the magnetic field in a cosmological simulation. It is often underestimated, because extremely high resolution is technically required to resolve the turbulent dynamo process in the intergalactic medium, which drives magnetic field amplification (e.g. Vazza et al. 2014, 2018; Böss et al. 2024; Steinwandel et al. 2024). Since the radio properties of the system are based on synchrotron emission, this can significantly alter the final result. To obtain more realistic predictions, it is therefore appropriate to compare the simulated magnetic field strength with analytic predictions for the system. In Fig. 7, we show the time-dependent radio luminosity due to synchrotron emission at for the simulated ORC. The three vertical lines represent the three times from the visualisation in Fig. 1, which are each about apart, starting shortly before the merger. The coloured regions represent three different models for the magnetic field strength, where the lower end is the prediction by the respective model and the upper end represents a boost by a factor of 10. The orange area shows the luminosity profile determined by the simulated magnetic field strength , while the other two were generated by assuming a balance between i) magnetic field pressure and thermal pressure (, blue) and ii) magnetic field pressure and turbulent pressure (, red). Hence, the local magnetic field strength is given by
@@
respectively in the two cases, where is the thermal pressure, the mass density and the turbulent velocity. The latter is defined as , where is the local velocity. The local bulk velocity is computed as the kernel weighted mean velocity of the neighbouring particles
@@
In order to keep the models realistic overall, the value of the plasma- value of 20 and the balance factor were scaled for the whole simulation such that the field strength in the central regions of the merging haloes is of the order of several , which is in accordance with measurements for galaxies in this mass range (e.g. Chyży & Beck 2004; Opher et al. 2009; Nikiel-Wroczyński et al. 2020). As expected, the two models generally produce higher values for the total magnetic field strength than the simulated one, reflected by the systematically greater inferred radio power. The turbulent model is overall the most powerful and is particularly visible at later times. At , i.e. when we identified the structure as an ORC (see Fig. 3), the maximum of the field strength distribution of the respective particles is (the histogram can be found in the Appendix in Fig. 11). Recent polarimetry measurements of ORC J0356–4216 (Taziaux et al. 2025) report field strengths of the same order of magnitude, which are about twice as large as in our simulated ORC.
@@
The results from the pilot study by Dolag et al. (2023) introduced in the beginning of this Section are visualised as the horizontal grey band in Fig. 7, where its width represents the overall extent of the total radio power inferred in post-processing of their simulation, ranging between and at (cf. Fig. 9 by Dolag et al. 2023). Since their observer frequency was at a higher value than in our figure555The total luminosity for observed objects is more often published at this lower frequency. Therefore we adopt this same frequency to enable comparison with more counterparts., we boost the quantities by Dolag et al. (2023) by a factor of ten666Such scaling for synchrotron emissivity is valid for a relatively hard energy spectrum, as in our case, and can be validated by comparing the power at and in Fig. 4. Compared to their estimate, our simulation exceeds it by at least three orders of magnitude at the moment of peak brightness – already when using the radio power inferred from the simulated magnetic field strength alone. But even taking the boost of stronger field models into account, we still do not reach the luminosity of observed counterparts indicated by the various dashed-dotted lines in Fig. 7. Consequently, also the surface brightness is under-predicted by the simulation. Physalis, for example, has a flux density of the order of in the outer radio rings ( shown in the Appendix in Fig. 12), while the simulated counterpart is about three orders of magnitude lower at (see Fig. 3). This means that, although we now resolve a halo in the correct mass regime and reproduce the expected X-ray properties (see Sect. 3.2), there is clearly still a missing aspect in the picture, that prevents us from fully explaining the physical nature of these systems.
@@ 4.2 Potential role of AGN and stellar feedback
Assuming no other essential process produces significant radio emission besides synchrotron emission, this lack of total radio luminosity implies insufficient relativistic CR electrons present and/or underestimated local magnetic field strengths. Since the maximum Mach number of the shock, when the simulated ORC is also the brightest, is already relatively large (, see Sect. 3.4), a stronger shock would not remedy the CR deficiency, because their acceleration efficiency via DSA would not increase substantially. A stronger magnetic field would in principle also give rise to stronger synchrotron emission. The problem is that several processes currently under-represented in our magnetic field models are presumed to contribute to the magnetic field state in the medium, for example Bell amplification in shocks or halo magnetisation by highly magnetised or laminar outflows. To test the effect of stronger magnetic fields in our simulation, we would need to increase the field strength across the whole system in order to stay self-consistent. But this would push the field strength in the central regions of the galaxies beyond actually observed values, and is therefore not an option. Additionally, we need to consider the relative importance of fossil CR populations in the halo, which would be present following prior injection events of CRs into the CGM of the merging galaxies. Typically, enough time would have passed for these old populations to cool down to mildly relativistic energies – meaning they would no longer be radio-bright – but not enough time for them to join the thermal pool of electrons. While a strong shock is necessary to accelerate electrons from the Maxwellian distribution, a relatively weak shock would already suffice to push such fossil CRs back to relativistic energies again, effectively boosting the total radio brightness of the whole system (e.g. Caprioli et al. 2018; Kang 2018).
Apart from merger or accretion shocks, stellar outflows and AGN feedback are the other main sources for these fossil populations, and it can be assumed that they should play a role in the two galaxies in our simulation. But since it was performed in a non-radiative fashion, meaning that it did not include star formation, it is not possible to make definite statements about the actual contribution of stars and black holes at this point. To our knowledge, no studies exist yet about the relative importance of shocks versus stellar and AGN contributions in the energy budget of the CGM. Shabala et al. (2024) give an expression to estimate the total energy injected into the medium by a pair of jets in their Eq. A6. Assuming that a violent AGN outburst can inflate a sphere with a radius of 200 kpc into an ambient medium with a density of , this would amount to a total energy budget from this AGN episode of . Studies of clusters with halo masses indicate that major mergers can release energies of the order of in these systems (Feretti et al. 2012; Dolag et al. 2023). We can scale this value down to group scales by assuming that this energy budget scales with the radio power versus halo mass relation, which persists across group and cluster masses. Thus, we arrive at an energy release of in groups with masses around (cf. Fig. 2 by Cuciti et al. 2023). Hence, according to this rudimentary estimate, this would mean that the energy releases by AGN and mergers are on the same order of magnitude inside a group environment.
@@
But even if AGN could contribute significantly more energy in relativistic electrons compared to merger events, it would not resolve the issue at hand, as becomes apparent from Fig. 7. There, the upper end of the inferred luminosity profiles from the simulation equates to an energy boost by a factor of 10, which evidently, would not suffice to reach the observed values. Nevertheless, the fractional polarisation actually also points towards the importance of fossil CRs in this system. The simulation produces values mainly between (see Sect. 3.5), while measurements of ORC 1 and ORC J0356–4216 report lower values mostly between in the interior and reaching 30% at most in the ring (Norris et al. 2022; Taziaux et al. 2025). If there are significant amounts of old CR populations present, one would certainly increase the number of emission centres along the line of sight compared to the case where electrons were accelerated from the thermal pool only. But since also the direction of the magnetic field vector varies along the line of sight—and the polarisation angle depends on that direction—this would naturally diversify the distribution of polarisation angles present in the total emission. Therefore, the total polarisation fraction along a given line of sight would decrease with additional fossil CRs populations, which is necessary to agree with observations. Thus, in order to provide a concluding answer to the question whether fossil CR populations are indeed the missing key to explaining the radio features of ORCs, one would need to perform the simulation again with star formation and proper feedback implementation, including self-consistent CR treatment. This remains a challenge for current state-of-the-art cosmological codes and will be addressed in future work.
4.3 Outlook on other formation scenarios
In addition to the merger shock paradigm, a variety of other formation pathways has been introduced in the literature so far. Among these, only two models have been tested via simulations as well—(i) the AGN driven model and (ii) the starburst wind model.
The most recent study on the AGN model was conducted by Shabala et al. (2024), where the authors presented hydrodynamical simulations of an AGN jet, where the synchrotron emissivity was estimated in post-process. Shabala et al. (2024) argued that AGN induced material would rise and evolve too slowly in comparison with the energy loss timescale of embedded CRs, meaning that AGN ejections alone cannot explain the extent and luminosity of ORCs. In order to examine the feasibility of CR re-acceleration, they set up a shock to traverse through the existing AGN bubble and obtained promising results that produced correct morphologies. Also as an idealized hydrodynamical simulation, Coil et al. (2024) recently presented a study on the starburst wind model, where the authors analysed the shocks induced by a sudden wind within an isolated galaxy and reported shock morphologies extending over several hundred of kiloparsecs.
Since both these models produce correct sizes, the next step would be to quantify the total radio power that can be reached via these formation pathways and compare it to observed counterparts. In particular, it is imperative to obtain a polarisation estimate from magnetohydrodynamic simulations of both models, as this property can change significantly depending on the underlying physical processes—and as such might be essential in estimating the likelihood of different formation scenarios.
@@ 5 Summary and conclusion
In this article, we presented a first of its kind zoom-in simulation of a galaxy group performed in a magnetohydrodynamic setup with spectral CR treatment on-the-fly. The scientific aim was to test the limits of a strong contender for a formation channel for ORCs – a new class of radio objects whose origin is still a puzzle. The idea is that during a major merger event between two galaxies, the immense energy dissipation can release several strong bow shocks (i.e. merger shocks). If the projection is close to perpendicular to the merger axis, these shocks can form a giant ring, that closely resembles the edge-brightened radio features now identified as ORCs in observations.
This simulation incorporates two major improvements over the pilot study by Dolag et al. (2023), who introduced this formation path: i) resolving a halo ten times more massive, with a final virial mass of about , now in agreement with the group mass measured at the centre of ORCs, and ii) determining the CR population and local magnetic field direction self-consistently, rather than in post-processing. We showed that the simulation reproduces the X-ray surface brightness of a galaxy group in that mass range and demonstrated that the peculiar offset between the X-ray and radio-bright centres observed in ORC Physalis (Koribalski et al. 2024b) can be explained by the difference in gas density at the centres of the merging galaxies (Sect. 3.2). By providing mock observations of the radio surface brightness and spectral index based on working frequencies of LOFAR, GMRT and ASKAP, we demonstrated that the morphology and size agree well with observations (Sect. 3.3). The emission spectrum of the radio shells indicates a sonic Mach number of , which is comparable to strong shock events in massive galaxy clusters (Sect. 3.4). The simulation produces a fractional polarisation degree of across the radio ring (Sect. 3.5). The pressure excess map provided in Sect. 3.6 reveals the sharp shock boundaries responsible for the radio features.
Compared to observations, the simulation under-predicts the total radio luminosity and over-estimates its polarisation degree (Sect. 4). Both issues would be alleviated by factoring in additional CR sources beyond DSA, the only acceleration mechanism in the present simulation. Such mechanisms would be injection of cosmic rays via stellar or AGN feedback, and could substantially contribute to the total energy budget and boost the radio brightness when re-accelerated by the merger shock. In addition, the simulation might underestimate the local magnetic field strength at the shock, leading to a significant under-prediction of inferred synchrotron intensity. We will implement descriptions of fossil CRs and magnetic field amplification mechanisms in future work to investigate whether they could explain the observed radio power of ORCs.